OpenStudy (anonymous):

(person who helps gets medal) Find the equation of a line that passes through the point (4,2) that is perpendicular to the line y = x. Show your work.

OpenStudy (anonymous):

\[y = \frac{ 4 }{ 3 } X\]

OpenStudy (anonymous):

is that your answer?

OpenStudy (anonymous):

no the fraction did not appear when I copy and paste

OpenStudy (anonymous):

the 4/3 is suppose to be in the question

OpenStudy (anonymous):

so what would the slope be of a line that is perpendicular to y=3/4x?

OpenStudy (anonymous):

How do we find that?

OpenStudy (anonymous):

mmm im having trouble since its a fraction ~

OpenStudy (anonymous):

do i put it like this y2 - y1 -------- x2 - x1

OpenStudy (anonymous):

If one line has a slope of \(\frac{1}{2}\), then the line that is perpendicular to it will be the negative reciprocal \(-2\)

OpenStudy (anonymous):

nope to find the slope of a line that is perpendicular to that, you just do the opposite reciprocal of the slope of y=3/4x

OpenStudy (anonymous):

so i make it negative ? 3 / 4 - ?

OpenStudy (anonymous):

and reciprocal

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

Yes, so now we have our new slope. Now we must plug in the other values that we're given. We're told that the line runs through the point (4, 2) Equation of a Line: \[\large (y-y_1)=m(x-x_1)\]

OpenStudy (anonymous):

Plug in the point into \(y_1\) and \(x_1\) and our slope \(m\) that we just also found.

OpenStudy (anonymous):

\[(y - 2) = -\frac{ 3 }{ 4 }( x - 4)\]

OpenStudy (anonymous):

this ?

OpenStudy (anonymous):

Yep! X) Now simplify. So distribute the (-3/4) and then add 2 to both sides.

OpenStudy (anonymous):

how would u distribute the -3/4 ?

OpenStudy (anonymous):

@CShrix

OpenStudy (anonymous):

Let's look at 2(x+1) By distributing, we multiply each term by 2 so 2(x+1) = 2x+2

OpenStudy (anonymous):

sooo you multiply the fraction by x and 4 ?

OpenStudy (anonymous):

could you draw it out instead of saying it? i can under stand it so much better

OpenStudy (anonymous):

 We have to multiply x and (-4) by (-3/4)

OpenStudy (anonymous):

so it would be -3/4x - 12/3 ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(3/4)(4) is not (12/3)

OpenStudy (anonymous):

what is it then ~

OpenStudy (anonymous):

im super bad at fractions

OpenStudy (anonymous):

\[y-2=-\frac{3}{4}x+\frac{12}{4}\]

OpenStudy (anonymous):

thats the answer ?

OpenStudy (anonymous):

Actually it depends on which form they want. If they want it in `slope-intercept `form, then we must add both sides by 2 to isolate y. But if they want the answer in `point-slope` form, then our answer is just \(\large (y - 2) = -\frac{ 3 }{ 4 }( x - 4)\) If it doesn't specify, then I would just leave it in point-slope form, which we already have.

OpenStudy (anonymous):

ok so yes alright thank you ! :)

OpenStudy (anonymous):

You are welcome :)